# On the speed of coming down from infinity for (sub)critical branching processes with pairwise interactions

### v1

@@ 2.2 Second order fluctuations in the subcritical cooperative regime
Our first main result in this section shows that the convergence established in Theorem 1 (see (6)) also holds in the $\mathbb{L}_{p}$ sense, for $p\in(0,2]$. However, this result requires the following additional assumption,
$\textrm{there exists}\quad\theta>e^{128\sqrt{2}}\quad\textrm{such that }\quad\sum_{i\geq 1}\theta^{i}\pi_{i}<\infty\quad\textrm{ and }\quad\sum_{i\geq 1}\theta^{i}b_{i}<\infty.$ (${\bf H4}$)
It is important to emphasise that the previous assumption can be established for any value of $\theta>1$, albeit at the cost of introducing a dependency of $\theta$ on the values of $p$ in the context of convergence within $\mathbb{L}_{p}$.
Theorem 3.
Suppose that $\mathbf{m}_{\rm int}<0$ and (${\bf H4}$) hold. Then for $p\in(0,2]$, we have
$\lim_{t\downarrow 0}\mathbb{E}\Big{[}\sup_{s\leq t}\Big{|}-\mathbf{m}_{\rm int}s\,Z_{s}^{(\infty)}-1\Big{|}^{p}\Big{]}=0.$
The key ingredient in the proof of Theorem 3 is an enhanced estimate, improving upon the one presented in [3, Lemma 4.6] (see also [3, Theorem 3.2]). This approach may be of independent interest, as it has the potential to be applied to a broader setting considered in [3]. It is important to note that the result in Theorem 3 actually holds for any $p>0$, provided that (${\bf H4}$) is satisfied for sufficiently large $\theta$, (see Section 3.4 for further details.).
Our last main result investigates the second-order fluctuations of $-\mathbf{m}_{\rm int}t\,Z_{t}^{(\infty)}$ for small times. More precisely, we shall study the asymptotic behaviour in a functional sense, as $\varepsilon\downarrow 0$, of the process $(X_{t}^{(\varepsilon)})_{t\geq 0}$, for $\varepsilon>0$, defined by
$X_{0}^{(\varepsilon)}=0\quad\text{and}\quad X_{t}^{(\varepsilon)}=\varepsilon^{-\frac{1}{2}}\Big{(}-\mathbf{m}_{\rm int}\varepsilon t\,Z_{\varepsilon t}^{(\infty)}-1\Big{)},\quad\text{for}\quad t>0.$ (10)
@@
$X_{0}=0\quad\text{and}\quad X_{t}=\frac{\bm{\sigma}_{{\rm int}}}{t}\int_{0}^{t}u{\rm d}W_{u},\quad\text{for}\quad t>0,$
@@
Theorem 4 recovers the second-order asymptotic result of the number of blocks in Kingman’s coalescent (i.e., when $d=\rho=0$ and $b=0$) about its speed of coming down from infinity [29, Theorem 1.1]. The process $(X_{t})_{t\geq 0}$ is similar to the process in [29, Theorem 1.3] in the case of the number of blocks of a $\Lambda$-coalescent with a Kingman part.
2.3 Comments about our results and existing literature
The proofs of Theorems 1 and 2 build upon [3, Theorem 4.5] (see Sections 3.2 and 3.3 for further details). However, addressing these results requires overcoming several technical challenges, particularly in determining the precise speed at which the processes come down from infinity. A central aspect of our approach is the representation of BPI processes as solutions to stochastic differential equations (SDEs) with jumps, enabling the effective application of martingale methods and stochastic calculus.
A complementary perspective on the phenomenon of coming down from infinity for BPI processes is provided in [7, Theorem 6]. This result offers a refined perspective using a moment duality approach. However, [7, Theorem 6] does not address the rate at which BPI processes come down from infinity, leaving the question open for subcritical and critical cooperative BPI processes within this framework.
To establish the connection between our results and [7, Theorem 6], we briefly outline the main findings of that work. Specifically, [7, Theorem 6] shows that a BPI process comes down from infinity if
$\mathcal{R}_{0,\rho}^{c,b}(\theta;1)\coloneqq\int_{\theta}^{1}\frac{\mathcal{Q}_{0,\rho}^{c,b}(\theta;u)}{u\Phi_{c,b}(u)}{\rm d}u=\infty\qquad\textrm{and}\qquad\mathcal{I}_{d,\rho}^{c,b}(\theta;1)\coloneqq\int_{\theta}^{1}\int_{\theta}^{u}\frac{1}{x\Phi_{c,b}(x)\mathcal{Q}_{d,\rho}^{c,b}(x;u)}{\rm d}x{\rm d}u<\infty,$ (11)
for some $\theta\in(0,1)$, where
$\mathcal{Q}_{d,\rho}^{c,b}(x;u)\coloneqq\exp\left\{\int_{x}^{u}\frac{\Psi_{d,\rho}(w)}{w\Phi_{c,b}(w)}{\rm d}w\right\}$ (12)
and for $u\in[0,1]$,
$\Psi_{d,\rho}(u)\coloneqq d-(\rho+d)u+\sum_{i\geq 1}\pi_{i}u^{i+1}\qquad\text{and}\qquad\Phi_{c,b}(u)\coloneqq c-(c+b)u+\sum_{i\geq 1}b_{i}u^{i+1}.$ (13)
In the subcritical cooperative regime, these conditions hold under the assumption that $\mathbf{m}_{\rm br}<\infty$, implying that the process does not explode; see [7, Example 1]. Additionally, in this case, we always have $\mathcal{I}_{d,\rho}^{c,b}(\theta;1)<\infty$, indicating that the associated BPI process comes down from infinity; see the comment following [7, Theorem 6].
In the critical cooperative regime, the situation is more nuanced. Lemmas 23 and 24 in Appendix A show that if conditions (${\bf H2}$) and (${\bf H3}$) are satisfied, then $\mathcal{R}_{0,\rho}^{c,b}(\theta;1)=\infty$ and $\mathcal{I}_{d,\rho}^{c,b}(\theta;1)<\infty$, ensuring that the process comes down from infinity. However, when (${\bf H2}$) holds and $\bm{\sigma}^{2}_{\rm int}$ is finite, we find that $\mathcal{I}_{d,\rho}^{c,b}(\theta;1)=\infty$, leaving us unable to determine whether the associated BPI process comes down from infinity. Notably, Theorem 2 (i) conclusively shows that a critical cooperative BPI process with finite $\bm{\sigma}^{2}_{\rm int}$ does not come down from infinity.
We now focus on Theorems 3 and 4. As previously noted, in assumption (${\bf H4}$), the condition $\theta>e^{128\sqrt{2}}$ can be relaxed to $\theta>1$. However, this adjustment comes at a cost: under our methodology, the $\mathbb{L}_{p}$ convergence in Theorem 3 will only hold for $p\in(0,\log(\theta)/\sqrt{8192}]$. Furthermore, if $\theta\leq e^{128\sqrt{2}}$, our approach is insufficient to establish the conclusions of Theorem 4. Assumption (${\bf H4}$) is thus critical to the robustness of our result in Theorem 4, and any relaxation of its conditions requires a thorough examination of its implications on convergence and related outcomes. Nonetheless, we conjecture that Theorems 3 and 4 may remain valid under weaker assumptions, warranting further investigation into alternative methodologies. In particular, we suspect that imposing moment conditions on the branching terms may not be necessary, as the dominant contributions arise from interactions whose rate of occurrence is quadratic.
The proof of Theorem 3 relies on a key improvement to the estimate provided in [3, Lemma 4.6], as detailed in Lemma 13. This refinement enables the application of a supermartingale inequality (see Lemma 22 in Appendix A), which closely resembles the approach used in the proof of [6, Theorem 2] where the $\mathbb{L}_{p}$ convergence for the speed of coming down from infinity for $\Lambda$ coalescent is studied.
A crucial component of the proof of Theorem 4 is the result in Theorem 3 for $p=2$. The proof of Theorem 4 builds upon the robust framework developed in [29, 30], where the fluctuations for the speed of coming down from infinity for $\Lambda$ coalescent are studied, with key modifications to account for the branching and cooperative dynamics of the BPI process.
@@ 2.4 Examples from population genetics
Kingman’s coalescent [26] is widely regarded as the simplest stochastic coalescence process. Despite its simplicity, it has profound mathematical properties that make it a cornerstone in population genetics. Informally, Kingman’s coalescent models the genealogical structure of large populations. As noted in the introduction (Section 1), the block counting process of Kingman?s coalescent corresponds to a subcritical cooperative BPI process with parameters $c>0$ and $d=\rho=b=0$. In particular, In particular, it satisfies $\mathbf{m}_{\rm int}<0$, $\mathbf{m}_{\rm br}<\infty$, (${\bf H1}$) and (${\bf H4}$), making the main results in Section 2 applicable.
@@
The Ancestral Selection Graph (ASG), introduced in [27, 31], extends Kingman’s coalescent by incorporating the effects of natural selection and mutation. Mutation modifies the genetic type of an ancestral line, potentially leading to its loss when tracing lineages backward in time. Selection, on the other hand, provides an evolutionary advantage to fitter individuals. Unlike the tree-like structure of Kingman’s coalescent, the ASG takes the form of a graph due to uncertainty introduced when tracing lineages backward without knowledge of individual types. Specifically, it is unclear whether an offspring results from standard reproduction or involves a fit individual. To capture this ambiguity, the ASG allows lineages to split into two potential paths. The block counting process of the ASG is a subcritical cooperative BPI process with parameters $c,d>0$, $b=0$, $\pi_{1}=\rho>0$ and $\pi_{i}=0$, for all $i\geq 2$. Here, $d$ accounts for lineage loss due to mutations, while $\rho$ captures lineage splitting caused by selection. Notably, $\mathbf{m}_{\rm int}<0$, $\mathbf{m}_{\rm br}<\infty$, and (${\bf H1}$) are satisfied, ensuring that the block counting process of the ASG comes down from infinity at the same speed as Kingman’s coalescent (Theorem 1); specifically, $\mathbf{m}_{\rm int}=-c$. Furthermore, (${\bf H4}$) is satisfied, and Theorem 4 applies with $\bm{\sigma}_{{\rm int}}^{2}=c$. These results were established for this specific model in [17].
@@
The dynamics of the block counting process in the ASG align with the so-called $(1,\rho,c,d)$-braco-process [16], which tracks the number of particles in a branching-coalescing system with the following rules: particles split into two at rate $\rho$, die at rate $d$, and any ordered pair of particles coalesces into one at rate $c$. Moreover, as highlighted in [17], the block counting process of the ASG bears similarities to the Ancestral Recombination Graph (ARG) [2, 1], a BPI-process characterised by parameters $c>0$, $b=d=0$, $\pi_{1}=\rho>0$ and $\pi_{i}=0$, for all $i\geq 2$.
@@
The Ancestral Selection/Efficiency Graph (ASEG), introduced in [15], generalises the ASG by incorporating the concept of efficiency, an ecological factor related to resource consumption strategies. Efficient individuals require fewer resources to produce progeny, giving them an advantage when resources are scarce, while inefficient individuals require more. The vertex counting process of the ASEG, as studied in [15], corresponds to a BPI process with parameters $c>0$, $b\in[0,c]$, $d=0$, $\pi_{1}=\rho\geq 0$ and $\pi_{i}=0$, for all $i\geq 2$. Here, $b$ represents the efficiency parameter, $\rho$ reflects selection, and there are no mutations $d=0$. In [15], the authors adopt the parametrisation $c=1/2$, $b=\kappa/2$, and $\rho=\alpha$, for $\kappa\in[0,1]$ and $\alpha\geq 0$.
When $b\in[0,c)$, the process is subcritical cooperative (i.e. $\mathbf{m}_{\rm int}<0$), satisfying $\mathbf{m}_{\rm br}<\infty$ and (${\bf H1}$). Thus, the vertex counting process of the ASEG comes down from infinity at the same speed (up to a constant) as Kingman’s coalescent (Theorem 1), with $\mathbf{m}_{\rm int}=-c+b$. Additionally, (${\bf H4}$) is satisfied, and Theorem 4 applies with $\bm{\sigma}_{{\rm int}}^{2}=c+b$. In the critical case where $c=b$, we have $\mathbf{m}_{\rm int}=0$ with $c>0$ and $\bm{\sigma}_{{\rm int}}^{2}<\infty$, and if $2\rho<\bm{\sigma}_{{\rm int}}^{2}$, assumption (${\bf H2}$) is satisfied. Consequently, Theorem 2(i) implies that the vertex counting process of the ASEG does not come down from infinity. However, when $2\rho\geq\bm{\sigma}_{{\rm int}}^{2}$, our result in the critical cooperative regime does not apply, as (${\bf H2}$) is not satisfied.
@@ 3 Proofs of the main results
Before we proceed with the proofs of our main result, we fix some notation. For two real-valued functions $f,g:[0,\infty)\rightarrow\mathbb{R}$, we write $f(x)\sim g(x)$, as $x\rightarrow y$, if $\lim_{x\rightarrow y}f(x)/g(x)=1$, where $y\in[0,\infty]$. We denote by $\overline{B}(x,\varepsilon)\coloneqq\inf\{y\in\mathbb{R}_{+}:|x-y|\leq\varepsilon\}$ the Euclidean closed ball centred at $x\in\mathbb{R}_{+}$ with radius $\varepsilon>0$. More generally, we denote by $\overline{B}_{d}(x,\varepsilon)\coloneqq\inf\{y\in\mathbb{R}_{+}:d(x,y)\leq\varepsilon\}$ the closed ball centred at $x\in\mathbb{R}_{+}$ with radius $\varepsilon>0$ associated with the application $d:\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$. Let $\Delta X_{s}=X_{s}-X_{s-}$ denote the jump at time $s\geq 0$ of a càdlàg process $(X_{t})_{t\geq 0}$.
@@ 3.1 Integral representation of BPI processes
$\nu_{\rm br}({\rm d}x)=d\delta_{-1}({\rm d}x)+\sum_{i=1}^{\infty}\pi_{i}\delta_{i}({\rm d}x)\qquad\text{and}\qquad\nu_{\rm int}({\rm d}x)=2c\delta_{-1}({\rm d}x)+\sum_{i=1}^{\infty}2b_{i}\delta_{i}({\rm d}x),$ (14)
@@
Set $\Delta\coloneqq\{(i,j)\in\mathbb{N}\times\mathbb{N}:1\leq i<j\}$. Let $N_{\rm br}$ and $N_{\rm int}$ be two independent $(\mathcal{F}_{t})_{t\geq 0}$-Poisson point measures on +××$\mathbb{R}_{+}\times\mathbb{N}\times\mathbb{N}_{-1}$ and +×$\mathbb{R}_{+}\times\Delta\times\mathbb{N}_{-1}$ with intensity measures ${\rm d}s\otimes\sum_{i\geq 1}\delta_{i}({\rm d}y)\otimes\nu_{\rm br}({\rm d}x)$ and ${\rm d}s\otimes\sum_{(i,j)\in\Delta}\delta_{(i,j)}({\rm d}y)\otimes\nu_{\rm int}({\rm d}x)$, respectively; where ${\rm d}s$ denotes the Lebesgue measure on $\mathbb{R}_{+}$. For $k\geq 2$ (an integer), let us denote $\Delta_{k}\coloneqq\{(i,j)\in\Delta:1\leq i<j\leq k\}$ and set $\Delta_{1}=\Delta_{0}=\emptyset$. Observe that $\nu_{\rm br}$ and $\nu_{\rm int}$ satisfy
$\int_{\mathbb{N}_{-1}}(1\wedge x^{2})\nu_{\rm br}({\rm d}x)<\infty\qquad\text{and}\qquad\int_{\mathbb{N}_{-1}}(1\wedge x^{2})\nu_{\rm int}({\rm d}x)<\infty,$ (15)
@@
$N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x)=\delta_{1}({\rm d}u)N_{\rm br}({\rm d}s,{\rm d}y,{\rm d}x)+\delta_{2}({\rm d}u)N_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x)$ (16)
be a Poison point measure on +×$\mathbb{R}_{+}\times\mathcal{X}$ with intensity measure
${\rm d}s\otimes q({\rm d}u,{\rm d}y,{\rm d}x)={\rm d}s\otimes\Big{(}\delta_{1}({\rm d}u)\otimes\sum_{i\geq 1}\delta_{i}({\rm d}y)\otimes\nu_{\rm br}({\rm d}x)+\delta_{2}({\rm d}u)\otimes\sum_{(i,j)\in\Delta}\delta_{(i,j)}({\rm d}y)\otimes\nu_{\rm int}({\rm d}x)\Big{)}.$ (17)
Proposition 1.
For $z\in\mathbb{N}_{0}$, the stochastic differential equation (SDE for short)
$Z_{t}=z+\int_{0}^{t}\int_{\mathcal{X}}H(Z_{s-},u,y,x)N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x),\quad t\geq 0,$ (18)
@@
$H(z,u,y,x)=\mathbf{1}_{\{u=1\}}x\mathbf{1}_{\{y\in\mathbb{N},0<y\leq z\}}+\mathbf{1}_{\{u=2\}}x\mathbf{1}_{\{y\in\Delta_{z}\}},\quad(z,u,y,x)\in\mathbb{N}_{0}\times\mathcal{X},$ (19)
has a unique strong solution. The solution $(Z_{t})_{t\geq 0}$ is a $\mathbb{N}_{0,\infty}$-valued $(\mathcal{F}_{t})_{t\geq 0}$-adapted process with càdlàg paths such that $\mathbb{P}(Z_{0}=z)=1$ and it satisfies (18) up to time $\tau_{n}\coloneqq\inf\{t\geq 0:Z_{t}\geq n\}$ for all $n\in\mathbb{N}$, and $Z_{t}=\infty$ for all $t\geq\tau\coloneqq\lim_{n\rightarrow\infty}\tau_{n}$. Moreover, the solution is a BPI-process issued from $z$.
@@
The proof follows from [32, Proposition 1] by considering the functions $b\equiv 0$ (no drift term), $\sigma\equiv 0$ (no Gaussian term), $h\equiv 0$ (no compensated Poisson measure term), $M=N$, $U=\mathcal{X}$ and $g=H$. The conditions of [32, Proposition 1] are verified using (15). ∎
We observe that the proof of Proposition 1 does not require any assumptions on the parameters of the BPI process. Henceforth, let $(Z_{t}^{(z)})_{t\geq 0}$ denote the BPI process issued from $z\in\mathbb{N}_{0}$, that is, the unique strong solution of the SDE (18). Note that the SDE representation (18) of the BPI-process bears some similarity to the construction of the standard Kingman coalescent, as explained in [29, Section 2]. Recall that the counting block process of the Kingman coalescent is a special case of a BPI process, when $c>0$, $d=\rho=b=0$.
Proposition 2.

### v2

@@ 2.2 Second order fluctuations in the subcritical cooperative regime
Our final main result investigates the second-order fluctuations of $\mathbf{m}_{\rm int}t\,Z_{t}^{(\infty)}$ for small times. More precisely, we will study the asymptotic behaviour in a functional sense, as $\varepsilon\downarrow 0$, of the process $(X_{t}^{(\varepsilon)})_{t\geq 0}$, for $\varepsilon>0$, defined by
$X_{0}^{(\varepsilon)}=0\quad\text{and}\quad X_{t}^{(\varepsilon)}=\varepsilon^{-\frac{1}{2}}\Big{(}-\mathbf{m}_{\rm int}\varepsilon t\,Z_{\varepsilon t}^{(\infty)}-1\Big{)},\quad\text{for}\quad t>0.$ (2.5)
@@
$X_{0}=0\quad\text{and}\quad X_{t}=\frac{\bm{\sigma}_{{\rm int}}}{t}\int_{0}^{t}u{\rm d}W_{u},\quad\text{for}\quad t>0,$ (2.6)
@@
Theorem 2.3 recovers the second-order asymptotic result of the number of blocks in Kingman’s coalescent (i.e., when $d=\rho=0$ and $b=0$) about its speed of coming down from infinity (Theorem 1.1 in [28]). The process $(X_{t})_{t\geq 0}$ is similar to the process in Theorem 1.3 of [28] in the case of the number of blocks of a $\Lambda$-coalescent with a Kingman part.
@@ 2.3 Examples from population genetics
Kingman’s coalescent [25] is widely regarded as the simplest stochastic coalescence process. Despite its simplicity, it has profound mathematical properties that make it a cornerstone in population genetics. As noted in the introduction (Section 1), the block counting process of Kingman’s coalescent corresponds to a subcritical cooperative BPI process with parameters $c>0$ and $d=\rho=b=0$. In particular, it satisfies $\mathbf{m}_{\rm int}<0$, $\mathbf{m}_{\rm br}<\infty$, ($\bf H1$}) and ($\bf H4$}), making the main results in Section 2 applicable.
@@
The Ancestral Selection Graph (ASG), introduced in [26, 30], extends Kingman’s coalescent by incorporating the effects of natural selection and mutation. Mutation modifies the genetic type of an ancestral line, potentially leading to its loss when tracing lineages backward in time. Selection, on the other hand, provides an evolutionary advantage to fitter individuals. Unlike the tree-like structure of Kingman’s coalescent, the ASG takes the form of a graph due to uncertainty introduced when tracing lineages backward without knowledge of individual types. Specifically, it is unclear whether an offspring results from standard reproduction or involves a fit individual. To capture this ambiguity, the ASG allows lineages to split into two potential paths. The block counting process of the ASG is a subcritical cooperative BPI process with parameters $c,d>0$, $b=0$, $\pi_{1}=\rho>0$ and $\pi_{i}=0$, for all $i\geq 2$. Here, $d$ accounts for lineage loss due to mutations, while $\rho$ captures lineage splitting caused by selection. Clearly, $\mathbf{m}_{\rm int}<0$, $\mathbf{m}_{\rm br}<\infty$, and ($\bf H1$}) are satisfied, ensuring that the block counting process of the ASG comes down from infinity at the same speed as Kingman’s coalescent (Theorem 2.1); specifically, $\mathbf{m}_{\rm int}=-c$. Furthermore, ($\bf H4$}) is satisfied, and Theorem 2.3 applies with $\bm{\sigma}_{{\rm int}}^{2}=c$. These results were established for this specific model in [16] (unfortunately, unpublished).
@@
The dynamics of the block counting process in the ASG align with the so-called $(1,\rho,c,d)$-braco-process [15], which tracks the number of particles in a branching-coalescing system with the following rules: particles split into two at rate $\rho$, die at rate $d$, and any ordered pair of particles coalesces into one at rate $c$. Moreover, as highlighted in [16], the block counting process of the ASG bears similarities to the Ancestral Recombination Graph (ARG) [2, 1], a BPI process characterised by parameters $c>0$, $b=d=0$, $\pi_{1}=\rho>0$ and $\pi_{i}=0$, for all $i\geq 2$.
@@
The Ancestral Selection/Efficiency Graph (ASEG), introduced in [14], generalises the ASG by incorporating the concept of efficiency, an ecological factor related to resource consumption strategies. Efficient individuals require fewer resources to produce progeny, giving them an advantage when resources are scarce, while inefficient individuals require more. The vertex counting process of the ASEG, as studied in [14], corresponds to a BPI process with parameters $c>0$, $b\in[0,c]$, $d=0$, $\pi_{1}=\rho\geq 0$ and $\pi_{i}=0$, for all $i\geq 2$. Here, $b$ represents the efficiency parameter, $\rho$ reflects selection, and there are no mutations $d=0$. In [14], the authors adopt the parametrisation $c=1/2$, $b=\kappa/2$, and $\rho=\alpha$, for $\kappa\in[0,1]$ and $\alpha\geq 0$. When $b\in[0,c)$, the process is subcritical cooperative (i.e. $\mathbf{m}_{\rm int}<0$), satisfying $\mathbf{m}_{\rm br}<\infty$ and ($\bf H1$}). Thus, the vertex counting process of the ASEG comes down from infinity at the same speed (up to a constant) as Kingman’s coalescent (Theorem 2.1), with $\mathbf{m}_{\rm int}=-c+b$. Additionally, ($\bf H4$}) is satisfied, and Theorem 2.3 applies with $\bm{\sigma}_{{\rm int}}^{2}=c+b$.
2.4 Comments about our results and organization of the rest of the paper
The proof of Theorem 2.1 builds upon Theorem 4.5 in [3] (see Section 3.2 for further details). However, addressing these results requires overcoming several technical challenges. Specifically, precisely determining the speed at which the processes come down from infinity. A central aspect of our approach is the representation of BPI processes as solutions to stochastic differential equations (SDEs) with jumps (see Section 3.1), allowing the effective application of martingale methods and stochastic calculus.
A complementary approach on the phenomenon of coming down from infinity for BPI processes is provided in Theorem 1.14 in [7]. This result offers a refined perspective using a moment duality approach. However, Theorem 1.14 in [7] does not address the rate at which subcritical (or critical) cooperative BPI processes come down from infinity, leaving the question open within this framework. In the subcritical cooperative regime, the conditions in [7] hold under the assumption that $\mathbf{m}_{\rm br}<\infty$, which implies that the process does not explode (see Example 2.1 in [7]). Additionally, in this case, the integral $\mathcal{I}_{d,\rho}^{c,b}(\theta;1)$ from [7] is always finite, which indicates that the associated BPI process comes down from infinity (see the comment following Theorem 1.14 in [7]).
In the critical cooperative regime, the situation is more delicate. There are situations where the test integral in Theorem 1.14 of [7] is unable to determine whether the associated BPI process comes down from infinity. For example, in the case where $\mathbf{m}_{\rm int}=0$ and $\mathbf{m}_{\rm br}<\bm{\sigma}^{2}_{\rm int}/2<\infty$, the integral $\mathcal{I}_{d,\rho}^{c,b}(\theta;1)$ from [7] is infinite (since we do not deal with the critical cooperative case here, details are left to the interested reader). Furthermore, the results from [3], specifically Proposition 4.4 and Theorem 4.5, are not straightforward to apply. Their use requires verifying non-trivial conditions that depend on the choice of a suitable function, denoted as $F$. While $F(z)\coloneqq\log(1+z)$ works for the subcritical cooperative regime (see Section 3.2), it is unclear if a similar function exists or can be found for the critical cooperative regime. Thus, we leave the question of the speed at which critical cooperative BPI processes come down from infinity as an open problem for future work.
We now focus on Theorems 2.2 and 2.3. As previously noted, in assumption ($\bf H4$}), the condition $\vartheta>e^{128\sqrt{2}}$ can be relaxed to $\vartheta>1$. However, this adjustment comes at a cost: under our methodology, the $\mathbb{L}_{p}$ convergence in Theorem 2.2 will only hold for $p\in(0,\log(\vartheta)/\sqrt{8192}]$. Furthermore, if $\vartheta\leq e^{128\sqrt{2}}$, our approach is insufficient to establish the conclusions of Theorem 2.3. Assumption ($\bf H4$}) is thus critical to the robustness of our result in Theorem 2.3, and any relaxation of its conditions requires a thorough examination of its implications. Nonetheless, we conjecture that Theorems 2.2 and 2.3 may remain valid under weaker assumptions. This leaves open the possibility of further investigation. In particular, we suspect that imposing moment conditions on the branching parameters may not be necessary, as the dominant contributions arise from interactions whose rate of occurrence is quadratic.
The proof of Theorem 2.2 (see Section 3.3) relies on Lemma 3.10, which is a key improvement to the estimate provided provided in Lemma 4.6 of [3] (see also Theorem 3.2 of [3]). This refinement enables the application of a supermartingale inequality (see Lemma A.1 in Appendix A), which resembles the approach used in the proof of Theorem 2 in [6] where the $\mathbb{L}_{p}$ convergence for the speed of coming down from infinity for $\Lambda$ coalescents was studied. (This approach may be of independent interest, as it has the potential to be applied to a broader setting, such as that considered in [3].)
A crucial component of the proof of Theorem 2.3 is the result in Theorem 2.2 for $p=2$ (see Section 3.4 for further details). The proof of Theorem 2.3 builds upon the robust framework developed in [28, 29], where the fluctuations for the speed of coming down from infinity for $\Lambda$ coalescents are studied, with key modifications to account for the branching and cooperative dynamics of the BPI process.
@@ 3 Proofs of the main results
Before we proceed with the proofs of our main results, we fix some notation. For two real-valued functions $f,g:[0,\infty)\rightarrow\mathbb{R}$, we write $f(x)\sim g(x)$, as $x\rightarrow y$, if $\lim_{x\rightarrow y}f(x)/g(x)=1$, where $y\in[0,\infty]$. We denote by $\overline{B}(x,\varepsilon)\coloneqq\inf\{y\in\mathbb{R}_{+}:|x-y|\leq\varepsilon\}$ the Euclidean closed ball centred at $x\in\mathbb{R}_{+}$ with radius $\varepsilon>0$. More generally, we denote by $\overline{B}_{d}(x,\varepsilon)\coloneqq\inf\{y\in\mathbb{R}_{+}:d(x,y)\leq\varepsilon\}$ the closed ball centred at $x\in\mathbb{R}_{+}$ with radius $\varepsilon>0$ associated with the application $d:\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$. Let $\Delta X_{s}=X_{s}-X_{s-}$ denote the jump at time $s\geq 0$ of a càdlàg process $(X_{t})_{t\geq 0}$.
@@ 3.1 Integral representation of BPI processes
$\nu_{\rm br}({\rm d}x)=d\delta_{-1}({\rm d}x)+\sum_{i=1}^{\infty}\pi_{i}\delta_{i}({\rm d}x)\qquad\text{and}\qquad\nu_{\rm int}({\rm d}x)=2c\delta_{-1}({\rm d}x)+\sum_{i=1}^{\infty}2b_{i}\delta_{i}({\rm d}x),$ (3.1)
@@
Set $\Delta\coloneqq\{(i,j)\in\mathbb{N}\times\mathbb{N}:1\leq i<j\}$. Let $N_{\rm br}$ and $N_{\rm int}$ be two independent $(\mathcal{F}_{t})_{t\geq 0}$-Poisson point measures on $\mathbb{R}_{+}\times\mathbb{N}\times\mathbb{N}_{-1}$ and $\mathbb{R}_{+}\times\Delta\times\mathbb{N}_{-1}$ with intensity measures ${\rm d}s\otimes\sum_{i\geq 1}\delta_{i}({\rm d}y)\otimes\nu_{\rm br}({\rm d}x)$ and ${\rm d}s\otimes\sum_{(i,j)\in\Delta}\delta_{(i,j)}({\rm d}y)\otimes\nu_{\rm int}({\rm d}x)$, respectively; where ${\rm d}s$ denotes the Lebesgue measure on $\mathbb{R}_{+}$. For $k\geq 2$ (an integer), let us denote $\Delta_{k}\coloneqq\{(i,j)\in\Delta:1\leq i<j\leq k\}$ and set $\Delta_{1}=\Delta_{0}=\emptyset$. Observe that $\nu_{\rm br}$ and $\nu_{\rm int}$ satisfy
$\int_{\mathbb{N}_{-1}}(1\wedge x^{2})\nu_{\rm br}({\rm d}x)<\infty\qquad\text{and}\qquad\int_{\mathbb{N}_{-1}}(1\wedge x^{2})\nu_{\rm int}({\rm d}x)<\infty,$ (3.2)
@@
$N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x)=\delta_{1}({\rm d}u)N_{\rm br}({\rm d}s,{\rm d}y,{\rm d}x)+\delta_{2}({\rm d}u)N_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x)$ (3.3)
be a Poison point measure on $\mathbb{R}_{+}\times\mathcal{X}$ with intensity measure
${\rm d}s\otimes q({\rm d}u,{\rm d}y,{\rm d}x)={\rm d}s\otimes\Big{(}\delta_{1}({\rm d}u)\otimes\sum_{i\geq 1}\delta_{i}({\rm d}y)\otimes\nu_{\rm br}({\rm d}x)+\delta_{2}({\rm d}u)\otimes\sum_{(i,j)\in\Delta}\delta_{(i,j)}({\rm d}y)\otimes\nu_{\rm int}({\rm d}x)\Big{)}.$ (3.4)
Proposition 3.1.
For $z\in\mathbb{N}_{0}$, the stochastic differential equation
$Z_{t}=z+\int_{0}^{t}\int_{\mathcal{X}}H(Z_{s-},u,y,x)N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x),\quad t\geq 0,$ (3.5)
@@
$H(z,u,y,x)=\mathbf{1}_{\{u=1\}}x\mathbf{1}_{\{y\in\mathbb{N},0<y\leq z\}}+\mathbf{1}_{\{u=2\}}x\mathbf{1}_{\{y\in\Delta_{z}\}},\quad(z,u,y,x)\in\mathbb{N}_{0}\times\mathcal{X},$ (3.6)
has a unique strong solution. The solution $(Z_{t})_{t\geq 0}$ is a $\mathbb{N}_{0,\infty}$-valued $(\mathcal{F}_{t})_{t\geq 0}$-adapted process with càdlàg paths such that $\mathbb{P}(Z_{0}=z)=1$ and it satisfies (3.5) up to time $\tau_{n}\coloneqq\inf\{t\geq 0:Z_{t}\geq n\}$ for all $n\in\mathbb{N}$, and $Z_{t}=\infty$ for all $t\geq\tau\coloneqq\lim_{n\rightarrow\infty}\tau_{n}$. Moreover, the solution is a BPI-process issued from $z$.
@@
The result follows from Proposition 1 in [31] by considering the functions $b\equiv 0$ (no drift term), $\sigma\equiv 0$ (no Gaussian term), $h\equiv 0$ (no compensated Poisson measure term), $M=N$, $U=\mathcal{X}$ and $g=H$. The conditions of Proposition 1 in [31] are verified using (3.2). ∎
We observe that the proof of Proposition 3.1 does not require any assumptions on the parameters of the BPI process. Henceforth, let $(Z_{t}^{(z)})_{t\geq 0}$ denote the BPI process issued from $z\in\mathbb{N}_{0}$, that is, the unique strong solution of the SDE (3.5). Note that the SDE representation (3.5) of the BPI process bears some similarity to the construction of the standard Kingman coalescent, as explained in Section 2 of [28]. Recall that the counting block process of the Kingman coalescent is a special case of a BPI process, when $c>0$, $d=\rho=b=0$.
Proposition 3.2.